# Sufficiently Regularized Nonnegative Quartic Polynomials are Sum-of-Squares

### v1

@@ A.3 Global Optimality Conditions for $m_{3}$
Appendix B Supplement for Classes of Quartically Regularized Polynomial
B.1 Proof for Theorem 3.3
Proof.
Assume $T=0$, since $s^{*}$ is the global minimum, we have $\nabla m_{3}(s^{*})=0$, according to (5), we can write
$m_{3}(s^{*}+v)-m_{3}(s^{*})=\frac{1}{2}\omega\underbrace{\begin{bmatrix}0&0&0&0&\\ 0&\nabla^{2}m_{3}(s^{*})&\sigma s\mathbf{1}^{T}&0\\ 0&\sigma\mathbf{1}s^{T}&\frac{\sigma}{2}\mathbf{1}\mathbf{1}^{T}&0\\ 0&0&0&0\end{bmatrix}}_{Q_{b}}\omega$ (65)
where $\phi_{2}$ defined as in (2), $\mathbf{1}=[1,\dotsc,1]^{T}\in\mathbb{R}^{n}$ and $\mathbf{1}\mathbf{1}^{T}\in\mathbb{R}^{n\times n}$ is a matrix with all entries equal to 1s. Let $\omega\in\mathbb{R}^{\frac{1}{2}(n+2)(n+1)}$ be any vector, we write $\omega=[1,v,u,z]^{T}$ as defined in (3). For any $\omega$, or equivalently for any $(u,v,z)$, we have
$2\omega^{T}Q_{b}(v)\omega$ $=$ $\nabla^{2}m_{3}(s^{*})[v]^{2}+2\sigma v^{T}s^{*}\sum_{i=1}^{n}u_{i}+\frac{\sigma}{2}\sum_{i=1}^{n}u_{i}^{2}+\sigma\sum_{1\leq i<j\leq n}u_{i}u_{j}$ (66)
$=$ $\nabla^{2}m_{3}(s^{*})[v]^{2}+\frac{\sigma}{2}\bigg(\sum_{i=1}^{n}u_{i}+2v^{T}s^{*}\bigg)^{2}-2\sigma(v^{T}s^{*})^{2}$
$=$ $\big[\underbrace{\nabla^{2}m_{3}(s^{*})-2\sigma s^{*}{s^{*}}^{T}}_{\succeq 0}\big][v]^{2}+\frac{\sigma}{2}\bigg(\sum_{i=1}^{n}u_{i}+2v^{T}s^{*}\bigg)^{2}\geq 0$
where the last inequality uses the second-order global optimality condition for $s^{*}$ given in Lemma 3.3 (i.e., $H+\sigma\|s^{*}\|I_{n}\succeq 0$). Thus, we obtained that $Q_{b}\succeq 0$ and therefore $m_{3}(s^{*}+v)-m_{3}(s^{*})$ is SoS. ∎
Remark B.1.
The bound in (66) is tight, as we can set $v=-s^{*}$ and $\sum u_{i}=\|s^{*}\|^{2}$. When $n=1$, Theorem 3.3 becomes the result in Remark 3.1.
Appendix C Supplement for Numerical Results
C.1 Experiment with MGH Test Set
In Table 2, * indicates that the algorithm failed to make progress after 3000 iterations. The Iter. and Eval. for AR3+SoS and AR3+SoS Convex excludes the unsolved SoS problems—these columns refer specifically to function evaluations and derivative evaluations. Problems marked with * are excluded from the average count. $\dagger$} indicates that the algorithm timed out: $\|\nabla f\|$ was reduced below $10^{-2}$ but did not reach the prescribed outer-loop tolerance of $10^{-3}$.
Table 2: Comparison of AR3+SoS Convex, AR3+SoS (Algorithm 1), ARC, and AR3+ARC across 20 test problems.
Problem Dim AR3 + SoS C AR3 + SoS ARC AR3 + ARC
$n$ Iter. Eval. Iter. Eval. Iter. Eval. Iter. Eval.
1 2 2656 2305 2656 2305 2317 2317 2354 2311
2 2 96 84 96 84 70 70 58 57
3 2 29 12 27 10 26 13 31 12
4 2 * * * * 1372 1372 32 31
5 2 67 59 67 59 57 57 58 57
6 2 113 100 113 100 104 104 119 99
7 3 815 711 792 691 779 779 717 716
8 3 825 723 825 723 707 707 726 725
9 3 19 11 18 11 24 17 17 12
10 3 * * * * * * * *
11 3 51 41 56 47 29 28 50 40
12 3 26 26 26 26 24 24 26 25
13 4 231 207 231 207 209 209 208 206
14 4 93 84 90 81 53 40 119 63
15 4 230 200 22 22 24 22 23 22
16 4 121 121 121 121 91 91 92 85
17 5 335 288 335 288 3000 2988 173 148
18 6 16 16 17 17 13 13 17 16
19 11 167 155 152 140 96 83 831 800
20 6 3000 2605 3000 2605 3000 3000 3000 3000
Average 494 431 480 419 631 628 455 443
C.2 Supplementary Results for Perturbing $m_{\mathcal{E}}$
@@
C.3 Derivative and Tensor for Sensor–Localization model

### v2

@@ A.5 Global Optimality Conditions for $m_{3}$
Appendix B Auxiliary Lemmas and Definitions
B.1 Definitions and Preliminaries for SoS Convex
Let $\mathbb{R}[s]^{n\times n}$ be the real vector space of $n\times n$ real polynomial matrices, where each entry of such a matrix is a polynomial with real coefficients.
Definition B.1.
(SoS-convex, [4, Def. 2.4] [20]) $h(s)$ is SoS-convex if its Hessian $\nabla^{2}h(s)$ is an SoS-matrix. A symmetric polynomial matrix $\check{H}(s)\in\mathbb{R}[s]^{n\times n}$ is an SoS-matrix if there exists a polynomial matrix $\check{V}(s)\in\mathbb{R}[s]^{n_{1}\times n}$ for some $n_{1}\in\mathbb{N}$, such that $\check{H}(s)=\check{V}(s)^{T}\check{V}(s)$.
Clearly, if $h(s):\mathbb{R}^{n}\rightarrow\mathbb{R}$ is SoS-convex, then $h(s)$ is convex [4, 23]. Recently, Ahmadi et al. [1] proved that for a locally strongly convex quartically regularized polynomial ( $H\succeq\delta I_{n}\succ 0$), there exists a sufficiently large regularization parameter $\sigma$ such that the polynomial becomes SoS-convex (see Theorem B.1).
Theorem B.1.
[…]
Lemma B.1.
B.2 Matrix Inequalities and Norms
Clearly, $\|T\|_{F}\geq\|T\|$. To see this, for a symmetric third-order tensor $T\in\mathbb{R}^{n\times n\times n}$, define its cubic (spectral) norm as
By the Cauchy–Schwarz inequality for tensor inner products, we have
Hence,
$\boxed{\|T\|\leq\|T\|_{F}.}$
and therefore
$\max_{1\leq i\leq n}|T_{iii}|\leq\|T\|.$
Appendix C Alternative Representations and Interpretations
C.1 Integral Representation of the Model
In this section, we present an alternative representation of the shifted polynomial
$q(v)=m_{3}(s^{*}+v)-m_{3}(s^{*})$
based on results from [20], and show its equivalence to the expansion used in the main text.
It is shown in Lemma 8 of [20] that if $s^{*}$ satisfies $\nabla m_{3}(s^{*})=0$, then
$m_{3}(s^{*}+v)-m_{3}(s^{*})=v^{T}\left[\int_{0}^{1}\int_{0}^{t}\nabla^{2}m_{3}(s^{*}+rv)\,dr\,dt\right]v.$ (53)
Define
$\mathcal{I}(s^{*},v):=\int_{0}^{1}\int_{0}^{t}\nabla^{2}m_{3}(s^{*}+rv)\,dr\,dt.$
Then $s^{*}$ is a global minimizer if and only if
$\nabla m_{3}(s^{*})=0\quad\text{and}\quad\mathcal{I}(s^{*},v)\succeq 0\quad\text{for all }v\in\mathbb{R}^{n}.$
We now show that this representation coincides with the expansion used in Lemma 2.1. Using
$\nabla^{2}m_{3}(s)=H+T[s]+\sigma\big(\|s\|^{2}I_{n}+2ss^{T}\big),$
we compute
$\mathcal{I}(s^{*},v)$ $=\int_{0}^{1}\int_{0}^{t}\Big[H+T[s^{*}+rv]+\sigma\big(\|s^{*}+rv\|^{2}I_{n}+2(s^{*}+rv)(s^{*}+rv)^{T}\big)\Big]drdt$
$=\frac{1}{2}\nabla^{2}m_{3}(s^{*})+\frac{1}{6}T[v]+\frac{\sigma}{6}\Big(2{s^{*}}^{T}vI_{n}+4s^{*}v^{T}\Big)+\frac{\sigma}{12}\Big(\|v\|^{2}I_{n}+2vv^{T}\Big).$ (54)
Multiplying by $v^{T}$ and $v$, we obtain
$\mathcal{I}(s^{*},v)[v]^{2}=\frac{1}{2}\nabla^{2}m_{3}(s^{*})[v]^{2}+\frac{1}{6}T[v]^{3}+\sigma{s^{*}}^{T}v\|v\|^{2}+\frac{\sigma}{4}\|v\|^{4},$
which matches the expression in (1).
Appendix D Additional Examples and Extensions
D.1 Proof for the univariate case
Proof for Equation 17.
Let $q(v):=m_{3}(s^{*}+v)-m_{3}(s^{*})$, using (1), we have
Proof for Theorem 3.1.
Since $s^{*}$ is the global minimizer, by Lemma 3.1, (17) holds and
[…]
We deduce that $Q_{a}\succeq 0$, therefore $m_{3}(s^{*}+v)-m_{3}(s^{*})$ is SoS. ∎
D.2 Proof for Optimality Conditions for Schnabel model
Proof for Theorem 3.3.
Let $q_{\mathcal{S}}(v):=m_{\mathcal{S}}(s+v)-m_{\mathcal{S}}(s)$. For any vector $s,v\in\mathbb{R}^{n}$, let $q_{\mathcal{S}}(v):=m_{\mathcal{S}}(s+v)-m_{\mathcal{S}}(s)$. We have
where $\nabla^{2}m_{\mathcal{S}}(s)=H+\sum_{j=1}^{k}(a_{j}a_{j}^{T})(b_{j}^{T}s)+\sum_{j=1}^{k}3\sigma_{j}(a_{j}a_{j}^{T})(a_{j}^{T}s)^{2}.$ Using (55) and (22), we have
Proof for Theorem 3.4.
  1. 1.
  2. 2.
[…]
D.3 Global Minimization and SoS Bounds: Additional Experiments
@@
D.4 Derivative and Tensor for Sensor–Localization model
@@
D.5 Adaptive Third Order Regularization Method with SoS Taylor Models
All experiments were performed on an AMD Ryzen 5 7600 6-Core Processor (3.80 GHz) with 16 GB of RAM. We test three methods: AR3 + SoS, fminunc, ARC (Adaptive Regularization with Cubics [13]), and AR3 + ARC ( Third-order method with ARC as the inner solver [14]).
Unless otherwise stated, the parameters for all methods were set to $\rho_{1}=0.1$, $\rho_{2}=0.9$, $\eta=0.1$ and $\gamma_{1}=0.5$, and $\gamma_{2}=2$, $\delta=0.1$, $\sigma_{0}=100$. The inner solvers used for each method were mcm from [12, Algorithm 6.1] for ARC, the ARC algorithm itself for AR3 + ARC, and the MSDP/MSOL moment-SDP solver from GloptiPoly 3 [21] with SEDUMI for AR3 + SoS. The outer-loop termination tolerance was set to $\|\nabla f\|<10^{-3}$, and the maximum iteration count at 3000. If an inner iteration was used, its tolerance was fixed at $10^{-6}$ and the maximum inner iteration count, at 3000.
if Moment SDP solved successfully (status==1) then
  Extract one global optimal solution, $s_{k}$.
else
Step 4: Regularization parameter update.
if $\rho_{k}>\rho_{1}$ then
 Successful Iter. $x_{k+1}:=x_{k}+s_{k}$;
else if $\rho_{k}>\rho_{2}$ then
else if $\rho_{k}\leq\rho_{1}$ then
Repeat with $k:=k+1$.